OpenStudy (anonymous):

7a-3=3-2a solve for a

OpenStudy (anonymous):

Bring a's to one side and constants to the other.

OpenStudy (anonymous):

\[7a - 3 = 3 - 2a\]\[7a - 3 + 2a = 3 - 2a + 2a\]\[9a - 3 = 3\]\[9a - 3 + 3 = 3 + 3\]\[9a = 6\]\[\frac{9a}{9} = \frac{6}{9}\]\[a = \frac{6}{9}\]

OpenStudy (anonymous):

9a-3=3

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

BTW, 6/9 can be further simplified, but it looks good @ChristianGeek

OpenStudy (anonymous):

True, sorry...the final answer should be: \[a = \frac{ 2 }{ 3 }\]